solve the D.E dy/dx = (6x^5-2x+1)/(cosy+e^y)


1
Expert's answer
2022-03-16T11:39:59-0400

"\frac{dy}{dx}=\frac{6x^5-2x+1}{\cos y +e^y}\\
(\cos y +e^y)dy=(6x^5-2x+1)dx\\
\int(\cos y +e^y)dy=\int(6x^5-2x+1)dx\\
\sin y+e^y=x^6-x^2+x+C."